### v1

@@ 1.1 Background
Among major attacks that threaten deep learning systems (e.g., model extraction, model inversion (MI), poisoning, and adversarial attacks) [37], the MI attack stands out due to its ability to extract sensitive information from the training dataset and compromise user privacy [27]. MI attacks were proposed by Fredrikson et al. [22], which exploits the relationship between input data and model learning parameters to recover sensitive information. The impact of MI attacks goes beyond violations of personal privacy, undermining trust in machine/deep learning systems in critical applications such as biometrics, healthcare analytics, and financial modeling. For example, using MI attacks on face recognition systems, an adversary can approximate face features in the training dataset and reconstruct the face image, thereby undermining the security of biometric systems [154, 113, 96]. In healthcare systems, leaked sensitive patient data may violate ethical standards and lead to legal consequences [15]. Similarly, in financial systems, private transaction data or credit scores could be compromised by MI attacks [71]. In addition to privacy concerns, these attacks can erode users’ trust in machine/deep learning systems, as users may lose confidence in the security of their personal or sensitive information. A visualization example of the MI attack is shown in Figure 1.
Figure 1: A visualization example of the model inversion (MI) attack (adapted from [76]).
@@ 1.2 Existing Surveys on MI Attacks and Defenses
1.3 Motivation and Contributions of This Work
The existing surveys mentioned above have advanced the categorization of MI attacks and their impact. However, these reviews often lack a unifying framework that integrates the latest advances across various paradigms and diverse applications, while also failing to provide a comprehensive summary of emerging challenges and corresponding future research directions. Many studies narrowly focus on specific techniques, datasets, or environments, leaving critical gaps in understanding broader MI threats, especially in complex distributed environments such as FL. Our survey aims to address these limitations through a comprehensive and integrated analysis of MI attacks and defenses. By combining theoretical insights with practical considerations, this survey aims to advance research on MI defense and guide the development of robust privacy protection strategies. The main contributions of this research are summarized below.
@@ 1.3 Motivation and Contributions of This Work
1.4 Structure of the Paper
This paper explores MI attacks and defenses, starting with the fundamentals of MI attacks (Section 2) to establish some key definitions and concepts. A taxonomy of MI attacks is provided in Section 3, including gradient inversion, generative model-based, and optimisation-driven approaches. Section 4 reviews the effectiveness of MI defense mechanisms, such as gradient perturbation and differential privacy. The paper also presents MI-related evaluation metrics (Section 5) and datasets (Section 6), followed by challenges in MI studies and future research directions (Section 7). Concluding remarks are given in Section 8.
@@ 2.1.1 Deep Learning Models
$h^{(l)}=f\left(W^{(l)}\cdot h^{(l-1)}+b^{(l)}\right)$
@@
$\hat{y}=\sigma\left(f_{n}\circ f_{n-1}\circ\ldots\circ f_{1}(\mathbf{x})\right)$
@@
Model Parameters. Given a deep learning model, weight ($W$) represents the connection strength between neurons in adjacent layers and is updated iteratively to improve the model’s predictions; and $\nabla W$ represents the gradient, obtained by computing the derivative of the loss function $\ell$ with respect to weight $W$. Gradients indicate the direction and magnitude of change needed in the weights to reduce the loss [168], expressed by
$\nabla W=\frac{\partial\ell}{\partial W}$
@@ 2.1.2 MI Attacks on Deep Learning Models
$(\mathbf{x}^{\prime*},\mathbf{y}^{\prime*})=\arg\min_{\mathbf{x}^{\prime},\mathbf{y}^{\prime}}\|\nabla W^{\prime}-\nabla W\|^{2}=\arg\min_{\mathbf{x}^{\prime},\mathbf{y}^{\prime}}\left\|\frac{\partial\ell(F(\mathbf{x}^{\prime},W),\mathbf{y}^{\prime})}{\partial W}-\nabla W\right\|^{2}$
@@ 3.1 Gradient Inversion Attacks
Gradient inversion attacks attempt to reconstruct private data from the gradients shared during the training of machine/deep learning models, especially in distributed or FL environments. These attacks utilize the gradients exchanged between clients and a central server to recover sensitive input data.
@@ 3.5 Summary
Structured around a host of key factors, Table 1 presents a detailed comparison of gradient inversion attacks and generative model-based attacks.
@@ 3.7 MI on Different Data Types
MI attacks present serious privacy risks to a panoply of data types, including image, audio, text, and tabular data. The latest advances in MI attacks increase the vulnerability of these data types to MI. MI attacks leverage the learned parameters, features or outputs of deep learning models to reconstruct sensitive training data, thereby compromising the privacy of user information. As deep learning systems increasingly penetrate areas such as visual recognition, speech processing, natural language understanding, and analysis of structured data, the possibility of privacy breaches by malicious attacks is increasing.
@@ 4.8 Summary
Defense strategies against MI attacks have evolved into diverse approaches to dealing with privacy vulnerabilities while attempting a good balance between efficiency, practicality, and robustness. These approaches are organized into six categories (see Sections 4.1-4.6). Some insights are: feature perturbation techniques (e.g., Crafter and Soteria) distort sensitive data while maintaining accuracy; gradient pruning methods (e.g., PATROL) improve privacy and communication efficiency by reducing the gradient; gradient perturbation strategies (e.g., GradPrivacy) show promise in obfuscating data in collaborative learning processes; DP remains crucial, and innovations such as class-level DP extend privacy guarantees; encryption methods, including certificateless homomorphic encryption, secure aggregation, and perceptual hashing, provide strong data protection by encrypting shared information; and model-enhancing techniques such as PRECODE and ResSFL protect privacy through architectural enhancements that defend against attacks without sacrificing performance. Despite the progress, challenges remain in balancing privacy, utility, and computational costs, especially in confronting advanced threats [107].
@@ 5 Evaluation Metrics in MI Attacks and Defenses
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left(x_{i}-y_{i}\right)^{2}$
where $x_{i}$ is the predicted value and $y_{i}$ the ground truth value. Smaller MSE values indicate better performance.
@@
$\text{LPIPS}(\mathbf{x},\mathbf{y})=\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h,w}\|\phi_{l}^{h}(\mathbf{x})-\phi_{l}^{h}(\mathbf{y})\|^{2}$
@@
$\text{PSNR}=10\cdot\log_{10}\left(\frac{\text{MAX}^{2}}{\text{MSE}}\right)$
@@
$\text{SSIM}(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})}$
where $\mu_{x},\mu_{y}$ are mean intensities of images $x$ and $y$; $\sigma_{x}^{2},\sigma_{y}^{2}$ are variances of $x$ and $y$; $\sigma_{xy}$ is the covariance between $x$ and $y$; and $C_{1},C_{2}$ are stabilizing constants.
@@
$\text{FID}=\|\mu_{x}-\mu_{y}\|_{2}^{2}+\text{Tr}(\Sigma_{x}+\Sigma_{y}-2(\Sigma_{x}\Sigma_{y})^{1/2})$
@@
$\text{FSIM}=\frac{\sum_{p\in\Omega}\text{PC}_{m}(p)\cdot\text{S}(p)}{\sum_{p\in\Omega}\text{PC}_{m}(p)}$
@@
$\text{AVD}(\mathbf{v}_{\text{source}},\mathbf{v}_{\text{target}})=\||\nabla\mathbf{v}_{\text{source}}||\nabla\mathbf{v}_{\text{target}}|\|+\||\nabla^{2}\mathbf{v}_{\text{source}}|-|\nabla^{2}\mathbf{v}_{\text{target}}|\|$
where $\nabla\mathbf{v}=\frac{\partial\mathbf{v}}{\partial i}+\frac{\partial\mathbf{v}}{\partial j}$ is the spatial gradient; $\nabla^{2}\mathbf{v}=\frac{\partial^{2}\mathbf{v}}{\partial i^{2}}+\frac{\partial^{2}\mathbf{v}}{\partial j^{2}}$ is the second-order spatial gradient; and $\mathbf{v}_{\text{source}}$ and $\mathbf{v}_{\text{target}}$ are the source and target images, treated as 2D arrays of pixel values.
@@
$\text{RDLV}=\frac{\text{SSIM}(T_{i},I_{i})-\text{SSIM}(T_{i},P)}{\text{SSIM}(T_{i},P)}$

### v2

@@ 1 Introduction
Among major attacks that threaten deep learning systems (e.g., model extraction, model inversion (MI), poisoning, and adversarial attacks) [44], the MI attack stands out due to its ability to extract sensitive information from the training dataset and compromise user privacy [33]. MI attacks were proposed by Fredrikson et al. [27], which exploits the relationship between input data and model learning parameters to recover sensitive information. The impact of MI attacks goes beyond violations of personal privacy, undermining trust in machine/deep learning systems in critical applications such as biometrics, healthcare analytics, and financial modeling. For example, using MI attacks on face recognition systems, an adversary can approximate face features in the training dataset and reconstruct the face image, thereby undermining the security of biometric systems [171, 127, 108]. In healthcare systems, leaked sensitive patient data may violate ethical standards and lead to legal consequences [20, 129, 91]. Similarly, in financial systems, private transaction data or credit scores could be compromised by MI attacks [80]. In addition to privacy concerns, these attacks can erode users’ trust in machine/deep learning systems, as users may lose confidence in the security of their personal or sensitive information. A visualization example of the MI attack is shown in Figure 1 (adapted from [86]).
Figure 1: A visualization example of the model inversion (MI) attack.
@@ 1.1 Existing Surveys on MI Attacks and Defenses
Differences to Existing Surveys. The main differences between our work and existing surveys can be summarized as follows. First, although the existing surveys have advanced the categorization of MI attacks and their impact, these reviews lack a unifying framework that integrates the latest advances across various paradigms and applications. Second, they fail to provide a comprehensive summary of emerging challenges and corresponding future research directions. Third, many of these studies narrowly focus on specific techniques (e.g., only about gradient inversion attacks and defenses), leaving critical gaps in understanding broader MI attacks and defenses.
Our survey aims to address aforementioned limitations through a comprehensive and integrated analysis of MI attacks, defenses and other aspects (e.g., applications, data types, and evaluation metrics). By combining theoretical insights with practical considerations in different applications, our work seeks to advance research on MI defense and guide the development of robust privacy protection strategies. While almost all MI-related surveys cover some common topics such as attacks, defenses, and future research, our work provides a more in-depth review as it explores additional aspects that are not covered by existing surveys, such as different recommendations regarding future research directions. Moreover, we create a unified resource repository for studying MI attacks and defenses. This will benefit both novice and experienced researchers in the field of deep learning-related security and privacy. A comparison between existing surveys and our work is presented in Table 1.
Surveys Gradient Inversion Attacks Generative Model-based Attacks Different Applications Different Data Types Defenses Evaluation Metrics Datasets Challenges and Future Research Resource Repository
[178] Yes No No No Yes No No Yes No
[18] Yes Yes Yes No Yes No No Yes No
[162] Yes No No No No No No Yes No
[29] Yes Yes No Yes Yes No No Yes Yes
[71] Yes Yes No No Yes No No No No
[116] Yes No No No Yes No No Yes No
This Survey Yes Yes Yes Yes Yes Yes Yes Yes Yes
Table 1: A comparison of existing surveys and this work. In this table, “Yes” indicates that the topic is covered, while “No” means that it is not covered.
1.2 Contributions of This Work
The main contributions of this research are summarized below.
@@ 1.2 Contributions of This Work
1.3 Literature Selection Methodology
This study employs a systematic method to identify, analyze, and synthesize relevant research articles on deep learning model inversion attacks and defenses. A thorough search is conducted across multiple reputable digital libraries, including IEEE Xplore, ACM Digital Library, SpringerLink, ScienceDirect, arXiv, and top-tier conferences such as CVPR, NeurIPS, ICML, ECCV, USENIX, and NDSS. The literature search covers foundational studies as well as latest advancements in the areas of interest. Both journal and conference papers, as well as preprints and technical reports, are included so as to achieve a complete coverage and a comprehensive overview.
To retrieve relevant studies, Boolean search queries are formulated with primary keywords, including “Model Inversion Attack”, “Gradient Inversion Attack”, “Privacy Breach in Deep Learning”, “Federated Learning Security”, “Privacy-Preserving AI”, “Generative Model Inversion”, “GAN-based Model Inversion”, and “Data Reconstruction Attack”. The Population, Concept, and Context framework [130] is applied to refine the search scope. The population consists of deep learning architectures, including CNNs, RNNs, transformers, and FL models. The concept focuses on techniques and defenses related to MI attacks, while the context covers AI security, biometric recognition, healthcare privacy, and cloud-based AI services.
A structured filtering process is implemented using Zotero 7.0 for reference management and duplicate removal. The inclusion criteria ensure the selection of peer-reviewed journal and conference papers on MI attacks and defenses, studies introducing novel attack techniques such as GAN-based, black-box, and gradient inversion attacks, experimental research evaluating attack effectiveness and defense robustness, and review papers summarizing up-to-date developments on topics pertaining to deep learning MI attacks and defenses. The exclusion criteria eliminate studies lacking empirical validation or theoretical discussions without experimental results, redundant works, and short papers, editorials, and opinion articles without technical depth. As a result, about 500 research articles are retrieved from various sources. Following a careful assessment to titles, abstracts, and full-text review, it results in a final selection of approximately 180 high-quality papers for this survey. The selected articles are collated into four themes. The first theme is about MI attack methodologies. The second theme explores defensive countermeasures and examination methods. The third theme focuses on evaluation metrics and datasets for MI attack research. The fourth theme addresses emerging challenges and future research directions.
@@ 2 Fundamentals of MI Attacks
MI attacks pose a significant threat to the privacy and security of deep learning models by attempting to reconstruct sensitive input data from model parameters, outputs, or intermediate representations. Understanding the underlying mechanisms of MI attacks is crucial for designing robust defense strategies. This section provides a comprehensive overview of MI attacks, beginning with their definition and characteristics, followed by an analysis of how these attacks exploit different components of deep learning models under various adversarial knowledge scenarios.
@@ 2.1 Definition and Characteristics
This subsection covers the mathematical formulation of deep learning models, their optimization process, and MI attacks that extract sensitive data from model parameters, outputs, or representations.
@@ 2.1.1 Deep Learning Models
$h^{(l)}=f\left(W^{(l)}\cdot h^{(l-1)}+b^{(l)}\right)$ (1)
@@
$\hat{y}=\sigma\left(f_{n}\circ f_{n-1}\circ\ldots\circ f_{1}(\mathbf{x})\right)$ (2)
@@
Model Parameters. Given a deep learning model, weight ($W$) represents the connection strength between neurons in adjacent layers and is updated iteratively to improve the model’s predictions. The loss function $\ell$ quantifies the discrepancy between the model’s predictions and the actual target values, guiding the optimization process. Gradient $\nabla W$ can be expressed by
$\nabla W=\frac{\partial\ell}{\partial W}$ (3)
where $\nabla W$ is the gradient of the weight matrix $W$, indicating the direction and magnitude of change needed to minimize the loss function $\ell$ [185]. The term $\frac{\partial\ell}{\partial W}$ in Equation (3) denotes the partial derivative of the loss function $\ell$ with respect to the weight matrix $W$, capturing how small changes in $W$ impact the loss.
@@ 2.1.2 MI Attacks on Deep Learning Models
$(\mathbf{x}^{\prime*},\mathbf{y}^{\prime*})=\arg\min_{\mathbf{x}^{\prime},\mathbf{y}^{\prime}}\|\nabla W^{\prime}-\nabla W\|^{2}=\arg\min_{\mathbf{x}^{\prime},\mathbf{y}^{\prime}}\left\|\frac{\partial\ell(F(\mathbf{x}^{\prime},W),\mathbf{y}^{\prime})}{\partial W}-\nabla W\right\|^{2}$ (4)
@@ 3.1 Gradient Inversion Attacks
FL (federated learning) is a decentralized approach where model training occurs across multiple devices, exchanging only model parameters instead of raw data to enhance privacy. In contrast, centralized learning (CL) collects all data on a central server for training purposes, offering efficiency but increasing privacy and security risks. The differences between FL and CL are demonstrated in Figure 3 (adapted from [18]). FL is still vulnerable to gradient inversion attacks, which aim to reconstruct private data from shared gradients during training. In FL environments, attackers can exploit the gradients exchanged between clients and the central server to recover sensitive input data, undermining the privacy benefits of decentralization.
Figure 3: Illustration of the differences between centralized learning (left) and federated learning (right).
@@ 3.5 Summary
Figure 4 presents an illustrative example of gradient inversion attacks vs. generative model-based attacks. Moreover, structured around a host of key factors, Table 2 presents a detailed comparison of gradient inversion attacks and generative model-based attacks.
Figure 4: A demonstrative example of gradient inversion attacks vs. generative model-based attacks.
@@ 3.6 MI in Different Applications
MI attacks pose a significant privacy threat across various domains where deep learning models process sensitive data. These attacks exploit learned representations to reconstruct private inputs, resulting in privacy and data breaches in biometric recognition systems, healthcare, and financial services.
@@ 3.7 MI on Different Data Types
MI attacks across different application domains, including biometric systems, healthcare systems, and financial systems and each of these applications processes different data types, such as images, audio, and text/tabular data, which are susceptible to MI attacks, as demonstrated in Figure 5. The latest advances in MI attacks increase the vulnerability of these data types to MI attacks. MI attacks leverage the learned parameters, features or outputs of deep learning models to reconstruct sensitive training data, thereby compromising the privacy of user information. As deep learning systems increasingly penetrate areas such as visual recognition, speech processing, natural language understanding, and analysis of structured data, the possibility of privacy breaches by malicious attacks is increasing.
Figure 5: MI attacks across various applications where different data types are processed.
@@ 3.7.3 MI on Texts and Tabular Data
3.8 MI on Emerging Generative AI and Foundation Models
With the recent rise of generative AI and foundation models like diffusion models (e.g., Stable Diffusion, DALL-E) and Large Language Models (LLMs) such as GPT-4, increased concerns about MI attacks have arisen. In contrast to traditional deep learning architectures, these models produce highly complex outputs and are therefore particularly vulnerable to data reconstruction attacks [28]. Attacks, such as MI attacks, seek to reconstruct training data from model outputs, raising ethical and security concerns. MI attacks leverage the model’s learnt representations to restore sensitive training data, posing substantial privacy risks in applications involving text, images, and audios. It is worth noting that MI attacks present significant privacy and security threats to generative AI models, including diffusion models, LLMs, and other foundation models. Although privacy-preserving techniques continue to advance, many challenges remain.
3.8.1 MI on LLMs
As deep learning evolves, LLMs (e.g., OpenAI’s GPT, Anthropic’s Claude, and Meta’s Llama) have greatly enhanced the ability to efficiently process a variety of downstream NLP tasks and unify them into a generative pipeline. Nevertheless, unrestricted access to these models also brings about potentially malicious attacks, such as MI attacks [61].
Embedding techniques are able to transform textual data into rich, dense numerical representations that capture semantic and syntactic properties, and thus have become a cornerstone of LLM functionality [75]. Embedding privacy is critical because these embeds often contain sensitive information about user data. Embedding inversion attacks are a specific subset of MI attacks that focus on restoring the original input data  [61] from feature embeddings. For example, Pan et al.  [103] transforms a text sequence into a set of words to perform multi-label classification to predict multiple words for given embeddings. Morris et al. [85] considered the issue of language model inversion, showing that the next-token probabilities contain a large amount of information about the preceding text. The approach demonstrates for the first time that language model predictions are mostly reversible. In a number of cases, the authors are capable of recovering inputs that are similar to the original text, sometimes even recovering the input text completely.
Chen et al. [12] studied the security of LLMs from the perspective of multilingual embedding inversion. Specifically, the authors defined the problem of black-box multilingual and cross-lingual inversion attacks with a special focus on cross-domain scenarios. The results show that multilingual models are more vulnerable to inversion attacks than monolingual models. Dai et al. [19] discovered a realistic attack surface of LLMs: privacy leakage of training data in decentralized training, and proposed the first activation inversion attack (AIA). The AIA utilizes a public dataset to construct a shadow dataset consisting of text labels and corresponding activations. Using this dataset, an attack model can be trained so as to reconstruct the training data from the activations in decentralized training. The authors conducted experiments on various LLMs and publicly available datasets to show the vulnerability of decentralized training to AIA.
Shu et al. [117] identified MI attacks in a split learning framework for LLMs, highlighting the necessity of security defenses. They introduced mutual information entropy for the first time to understand the information propagation of transformer-based LLMs and evaluated the privacy attack performance of LLM blocks. The authors proposed a two-stage attack system where the first stage projects representations into the embedding space, and the second stage uses a generative model to recover text from the embeddings. This work highlights comprehensively the potential privacy risks when deploying personalized LLMs at the edge side. To reduce embedding inversion attacks, Liu et al. [75] proposed Embedding Guard (Eguard). Eguard employs transformer-based projection networks and textual mutual information optimization to protect embeddings while retaining the utility of LLMs. This approach greatly reduces privacy risks and protects more than 95% of the tokens from being inverted, while maintaining high performance consistent with the original embeddings in downstream tasks.
3.8.2 MI on Diffusion Models
Diffusion models are emerging as favorable models for generating exceptionally high resolution image data. Diffusion models generate higher quality samples and are much easier to scale and control than previous models such as GAN. As a result, they have rapidly become the de facto approach for producing high-resolution images, and large-scale models such as DALL-E have attracted considerable public interest [9].
However, diffusion models like stable diffusion are also subject to MI attacks that can extract training data from diffusion models. Chen et al. [15] provided a comprehensive review of recent advances in image inversion techniques, highlighting two main paradigms: GAN inversion and diffusion model inversion. In the context of diffusion model inversion, the authors explored training-free strategies, fine-tuning methods, and the design of additional trainable modules, and emphasized their unique advantages and limitations. Carlini et al. [9] demonstrated that diffusion models memorize individual images from training data and emit these images at generation time. By using a generate-and-filter pipeline, the authors pulled over a thousand training examples from state-of-the-art models, varying from photos of individual people to company trademark logos. They also trained hundreds of diffusion models in a variety of settings to analyze how different modeling and data decisions impact privacy. The findings suggest that diffusion models are much less private than previous generative models such as GANs.
Huang et al. [42] investigated the privacy leakage risk of gradient inversion attacks. The authors designed a two-stage fusion optimization scheme that uses the trained generative model itself as prior knowledge to constrain the inversion search (latent) space, followed by pixel-level fine-tuning. The results show that the proposed optimization scheme can reconstruct an image almost identical to the original image. Wu et al. [146] presented a practical gradient inversion method, namely Deep Generative Gradient Inversion, which utilizes the diffusion model’s prior knowledge to improve the reconstruction performance for high-resolution datasets and larger batches. In addition, in order to address the spatial variation problem caused by the pre-trained diffusion model, group consistency regular terms are developed to constrain the distance between the reconstructed and aligned images. Zhang et al. [191] proposed a means of directing the inversion process of a diffusion model toward a synthetic embedding of the core distribution. In addition, the authors introduced a spatial regularization method to balance the attention to the concepts being composed. This is devised as a post-training method that can be seamlessly integrated with other inversion approaches. Experimental results show that the proposed method is effective in mitigating the overfitting problem and generating more diverse and balanced combinations of concepts in synthetic images.
@@ 4.8 Summary
Defense strategies against MI attacks have evolved into diverse approaches to dealing with privacy vulnerabilities while attempting a good balance between efficiency, practicality, and robustness. These approaches are organized into six categories (see Sections 4.1-4.6). Some insights are: feature perturbation techniques (e.g., Crafter and Soteria) distort sensitive data while maintaining accuracy; gradient perturbation strategies (e.g., GradPrivacy) show promise in obfuscating data in collaborative learning processes; DP remains crucial, and innovations such as class-level DP extend privacy guarantees; encryption techniques, such as certificateless homomorphic encryption, secure aggregation, and perceptual hashing, provide robust data protection by securing shared information; and model-enhancing techniques such as PRECODE and ResSFL protect privacy through architectural enhancements that defend against attacks without sacrificing performance.
Despite recent progress, it is challenging to achieve a balanced trade-off between privacy, utility, and computational costs, especially to counter advanced threats [120] and address security concerns in IoT and edge computing environments. Most of the MI defense mechanisms discussed earlier, except for gradient pruning methods such as PATROL [22], DGP [155], Guardian [26], and DGC [64], fail to consider the trade-off between efficiency and privacy preservation in IoT and edge computing settings. These environments are often constrained by limited computational resources, making traditional, resource-intensive security defenses impractical. Gradient pruning methods selectively prune gradients to enhance privacy and communication efficiency while maintaining model performance. For instance, DGC [64] reduces gradient exchange by 99.9% and still preserves model accuracy. However, nearly all existing gradient pruning methods have yet to be tested on real IoT or edge computing devices. Therefore, there is a pressing need for practical pruning strategies and lightweight MI defense mechanisms (e.g., efficient encryption techniques) to mitigate vulnerabilities while ensuring a feasible trade-off between defense performance and cost effectiveness (e.g., computational efficiency) in IoT and edge computing environments.
@@ 5 Evaluation Metrics in MI Attacks and Defenses
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left(x_{i}-y_{i}\right)^{2}$ (5)
where $x_{i}$ represents the predicted value and $y_{i}$ the ground truth value. Smaller MSE values indicate better performance.
@@
$\text{LPIPS}(\mathbf{x},\mathbf{y})=\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h,w}\|\phi_{l}^{h}(\mathbf{x})-\phi_{l}^{h}(\mathbf{y})\|^{2}$ (6)
@@
$\text{PSNR}=10\cdot\log_{10}\left(\frac{\text{MAX}^{2}}{\text{MSE}}\right)$ (7)
@@
$\text{SSIM}(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})}$ (8)
where $\mu_{x},\mu_{y}$ are mean intensities of images $x$ and $y$, respectively; $\sigma_{x}^{2},\sigma_{y}^{2}$ are variances of images $x$ and $y$, respectively; $\sigma_{xy}$ is the covariance between images $x$ and $y$; and $C_{1},C_{2}$ are stabilizing constants.
@@
$\text{FID}=\|\mu_{x}-\mu_{y}\|_{2}^{2}+\text{Tr}(\Sigma_{x}+\Sigma_{y}-2(\Sigma_{x}\Sigma_{y})^{1/2})$ (9)
@@
$\text{FSIM}=\frac{\sum_{p\in\Omega}\text{PC}_{m}(p)\cdot\text{S}(p)}{\sum_{p\in\Omega}\text{PC}_{m}(p)}$ (10)
@@
$\text{AVD}(\mathbf{v}_{\text{source}},\mathbf{v}_{\text{target}})=\||\nabla\mathbf{v}_{\text{source}}||\nabla\mathbf{v}_{\text{target}}|\|+\||\nabla^{2}\mathbf{v}_{\text{source}}|-|\nabla^{2}\mathbf{v}_{\text{target}}|\|$ (11)
where $\nabla\mathbf{v}=\frac{\partial\mathbf{v}}{\partial i}+\frac{\partial\mathbf{v}}{\partial j}$ is the spatial gradient; $\nabla^{2}\mathbf{v}=\frac{\partial^{2}\mathbf{v}}{\partial i^{2}}+\frac{\partial^{2}\mathbf{v}}{\partial j^{2}}$ is the second-order spatial gradient; and $\mathbf{v}_{\text{source}}$ and $\mathbf{v}_{\text{target}}$ are source and target images, treated as 2D arrays of pixel values.
@@
$\text{RDLV}=\frac{\text{SSIM}(T_{i},I_{i})-\text{SSIM}(T_{i},P)}{\text{SSIM}(T_{i},P)}$ (12)